OpenStudy (anonymous):

the sum of the reciprocals of two consecutive even integers is 9/40. This can be represented by the equation shown. 1/x + 1/x+2 = 9/40 use the rational equation to determine the integers. show all work.

OpenStudy (anonymous):

@jim_thompson5910 could u help with this??

jimthompson5910 (jim_thompson5910):

multiply every fraction by 40x(x+2)

OpenStudy (tkhunny):

x and x+2 are NOT necessarily EVEN numbers. If you solve as you have it set up, it will only be luck that gives you even numbers.

OpenStudy (anonymous):

i would add and then "cross multiply" i guess it is a matter of preference

OpenStudy (anonymous):

well i did what @jim_thompson5910 said and multiplied each fraction by 40x(x+2) and I got 40x+80 and 80x^2+160x

OpenStudy (anonymous):

would this be right??? @satellite73 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

1/x times 40x(x+2) = 40x+80 1/(x+2) times 40x(x+2) = 40x 9/40 times 40x(x+2) = 9x(x+2) = 9x^2+18x

jimthompson5910 (jim_thompson5910):

1/x + 1/(x+2) = 9/40 becomes 40x+80 + 40x = 9x^2 + 18x

OpenStudy (anonymous):

@jim_thompson5910 okay I see.. so where do we go from there?

jimthompson5910 (jim_thompson5910):

get the equation into standard form, then use the quadratic formula

jimthompson5910 (jim_thompson5910):

standard form for a quadratic is ax^2 + bx + c = 0

OpenStudy (anonymous):

@jim_thompson5910 so we use the quadratic formula on the equation you just found??

jimthompson5910 (jim_thompson5910):

no you have to get it into the form ax^2 + bx + c = 0 first

OpenStudy (anonymous):

@jim_thompson5910 okay so this is all i got ---> 9x^2+40x

jimthompson5910 (jim_thompson5910):

40x+80 + 40x = 9x^2 + 18x 80x+80 = 9x^2 + 18x 0 = 9x^2 + 18x-80x-80 0 = 9x^2 - 62x - 80 9x^2 - 62x - 80 = 0

jimthompson5910 (jim_thompson5910):

now you can use the quadratic formula

jimthompson5910 (jim_thompson5910):

you can factor if you want, but it might take a while lol

OpenStudy (anonymous):

@jim_thompson5910 okay so i got x= 8, -10/9

jimthompson5910 (jim_thompson5910):

same here

jimthompson5910 (jim_thompson5910):

so if x = 8, then x+2 = 10

jimthompson5910 (jim_thompson5910):

1/8 + 1/10 = 9/40 5/40 + 4/40 = 9/40 9/40 = 9/40

jimthompson5910 (jim_thompson5910):

works out, I'll let you check the other one

OpenStudy (anonymous):

@jim_thompson5910 so the answer would be x=8 , and -10/9??

jimthompson5910 (jim_thompson5910):

oh they have to be integers, so x = -10/9 must be ignored

jimthompson5910 (jim_thompson5910):

so the two integers are 8 and 10

OpenStudy (anonymous):

@jim_thompson5910 okay , thank you so much!

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

@dan815 @Hero if u guys help me i will fan u and medal :P

OpenStudy (anonymous):

@ytrewqmiswi

OpenStudy (anonymous):

okie so i don't get how they got x=8 and x=10 in the end why did -10/9 change to 10?

OpenStudy (anonymous):

help guys?

ganeshie8 (ganeshie8):

x = 8 or -10/9 since -10/9 is not an integer, ignore it.

OpenStudy (anonymous):

thats the part i don't understand

ganeshie8 (ganeshie8):

what part, could you elaborate

OpenStudy (ytrewqmiswi):

helpful link - https://answers.yahoo.com/question/index?qid=20130513215652AAFLnSA

OpenStudy (anonymous):

OMG THANKS MAN here a medal

OpenStudy (anonymous):

oh and ur first fan :)

OpenStudy (ytrewqmiswi):

haha XD thanks :D

OpenStudy (anonymous):

:P